OpenStudy (anonymous):

What value of x makes the equation true? –(9 + (–17)) = –9 + x

OpenStudy (anonymous):

To solve this problem, First compute the left side to equal a single numer then add nine to both sides to get x by itself and you should have your answer

OpenStudy (anonymous):

1st, distribute the - sign (which means multiply by -1) on the left side: -9 - (-17) = -9 + x Negative # times negative # = positive # -9 + 17 = -9 + x Add -9 and 17 together from the left side of the equation: 8 = -9 + x To get rid of the -9 from the right, add 9 to both sides (because -9 + 9 = 0): 17 = x

OpenStudy (anonymous):

SO MY answer is 17 im confused ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

You're right @sarai5